# Spatiotemporal Multi-Camera Calibration using Freely Moving People

### v2

@@ II-B1 Multi-camera Calibration
A common approach for multi-camera calibration utilizes reference objects such as checkerboard pattern [19] and ArUco markers [2] to facilitate establishing correspondences across multiple images. These methods, however, can be cumbersome and require careful operations. Leveraging dynamic foreground objects in our daily lives, such as freely moving people in scenes, has been proposed as a potential alternative to such manual calibration [20, 4, 21, 7, 22]. Using human heights in motion, represented as vertical line segments perpendicular to the ground plane, has demonstrated the possibility for calibration [21, 7]. This approach facilitated the estimation of vanishing points [21] and provided closed-form solutions for the line segments [7]. Additionally, human keypoints obtained from 2D human pose estimation have served as the calibration pattern et al. [20]. Moreover, 3D human poses, derived from elevating 2D human poses, have been employed as reference targets for calibration [4, 23]. Lee et al. [4] introduced a linear calibration algorithm based on factorization using 3D oriented points extracted from 3D human poses.
The methods discussed so far assume that the correspondences between views are determined beforehand, which lacks the versatility required for dynamic, multi-person scenes. To overcome this challenge, various priors, such as geometry, appearance, and motion, have been explored [24, 5, 25, 26]. Huang et al. [24] jointly recovers multiple human motions and extrinsic camera parameters using physics-based constraints to denoise 2D poses of pre-established correspondences and ensure temporal coherence with a human motion prior. Yan et al. [5] employed a re-identification (re-ID) network to associate bounding boxes across different camera views, then converted these into point correspondences to calculate relative camera poses across camera pairs. Xu et al. [25] treated the cross-view matching problem as a clustering problem using 2D appearance features within uncalibrated camera networks. Li et al. [26] uses 3D Re-ID appearance features from RGBD images to establish correspondences, which facilitates depth-based camera pose and 3D multi-human pose estimation.
@@ V-A Experimental Setups
For our experiments, we uniformly subsample the sequences within each scenario and apply random initial rotations. Trackings are initiated from each of subsampled frames for $T$ frames. We set $T$ between 5 and 15 frames, depending on the scenes. In the following experiments with unsynchronized setups, the temporal offset is set to $+5$ frames relative to the first camera with a search range of temporal candidates of ±$\pm 10$ frames. The optimization begins with initial offsets set to zero.
@@
CMU Panoptic [28] ZJU-MoCap [40] MMPTRACK [41]
@@
Voet al. [11] $E_{R}$[$\mathrm{rad}\mathrm{]}$ 0.03 - 0.006 - - 0.013 - - -
@@ V-B Generalizability to Different Scenes
We validated our algorithm on three datasets: CMU Panoptic [28], ZJU-MoCap [40], and MMPTRACK dataset [41], which includes a variety of camera configurations and human activities across both indoor and outdoor environments, as shown in Fig. 3 and Fig  4. Table II provides a summary of these datasets. In each scenario, we selected four cameras, each observing two or more people interacting freely throughout a clip of 300 to 1000 frames. For the CMU Panoptic dataset, we used a set of HD cameras.

### v3

@@ II-B1 Multi-camera Calibration
A common approach for multi-camera calibration utilizes reference objects such as checkerboard pattern [19] and ArUco markers [2] to facilitate establishing correspondences across multiple images. These methods, however, can be cumbersome and require careful operations. Leveraging dynamic foreground objects in our daily lives, such as freely moving people in scenes, has been proposed as a potential alternative to such manual calibration [20, 4, 21, 7, 22]. Using human heights in motion, represented as vertical line segments perpendicular to the ground plane, has demonstrated the possibility for calibration [21, 7]. This approach facilitated the estimation of vanishing points [21] and provided closed-form solutions for the line segments [7]. Additionally, human keypoints obtained from 2D human pose estimation have served as the calibration pattern [20]. Moreover, 3D human poses, derived from elevating 2D human poses, have been employed as reference targets for calibration [4, 23]. Lee et al[4] introduced a linear calibration algorithm based on factorization using 3D oriented points extracted from 3D human poses.
The methods discussed so far assume that the correspondences between views are determined beforehand, which lacks the versatility required for dynamic, multi-person scenes. To overcome this challenge, various priors, such as geometry, appearance, and motion, have been explored [24, 5, 25, 26]. Huang et al[24] jointly recovers multiple human motions and extrinsic camera parameters using physics-based constraints to denoise 2D poses of pre-established correspondences and ensure temporal coherence with a human motion prior. Yan et al[5] employed a re-identification (re-ID) network to associate bounding boxes across different camera views, then converted these into point correspondences to calculate relative camera poses across camera pairs. Xu et al[25] treated the cross-view matching problem as a clustering problem using 2D appearance features within uncalibrated camera networks. Li et al. [26] uses 3D Re-ID appearance features from RGBD images to establish correspondences, which facilitates depth-based camera pose and 3D multi-human pose estimation.
@@ V-A Experimental Setups
For our experiments, we uniformly subsample the sequences within each scenario and apply random initial rotations. Trackings are initiated from each of subsampled frames for $T$ frames. We set $T$ between 5 and 15 frames, depending on the scenes. In the following experiments with unsynchronized setups, the temporal offset is set to +5+5+ 5 frames relative to the first camera with a search range of temporal candidates of ±10$\pm$ 10± 10 frames. The optimization begins with initial offsets set to zero.
@@
CMU Panoptic [28] ZJU-MoCap [42, 40] MMPTRACK [41]
@@
Voet al[11] $E_{R}$[$\mathrm{rad}\mathrm{]}$ 0.03 - 0.006 - - 0.013 - - -
@@ V-B Generalizability to Different Scenes
We validated our algorithm on three datasets: CMU Panoptic [28], ZJU-MoCap [42, 40], and MMPTRACK dataset [41], which includes a variety of camera configurations and human activities across both indoor and outdoor environments, as shown in Fig. 3 and Fig  4. Table II provides a summary of these datasets. In each scenario, we selected four cameras, each observing two or more people interacting freely throughout a clip of 300 to 1000 frames. For the CMU Panoptic dataset, we used a set of HD cameras.